# Correlated Systematic Uncertainties and Errors-on-Errors in Measurement Combinations with an Application to the $7$-$8$ TeV ATLAS-CMS Top Quark Mass Combination

### v3

@@ 5.1 GVM analysis for combined estimate of $m_{t}$
Figure 1 illustrates the variation in the central value (i.e., Maximum Likelihood Estimator) of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. Specifically, the plot shows the dependence of the central value $\hat{\mu}$ on the error-on-error parameter $\varepsilon_{s}$ for each systematic uncertainty. The nomenclature for the systematic components utilized here is the same used in bib:ATLAS/CMS , where a detailed explanation of what they describe can be found. The most important conclusion here is that the central value of the combination is quite robust to the presence of uncertain systematic errors. Specifically, the change in the central value remains always within $0.1$ GeV for the explored range of $\varepsilon_{s}$ values, which is well within the confidence interval of approximately $0.3$ GeV. Ultimately, all central values converge to the BLUE result as $\varepsilon_{s}$ approaches zero, meaning that the way we implement correlations in our model, as described in Sec. 2, is consistent. The reason why the b tagging and QCD radiation lines do not converge to the same point is due to the regularization of their associated correlation matrices, as previously described.
@@
Figure 2 confirms our previous conclusion: the combination is generally robust to the presence of uncertain systematic errors. However, an increase of about $10\%$ in the confidence interval is observed when the b-JES systematic uncertainty is known with a relative error of approximately $50\%$. The combination’s heightened sensitivity to uncertainties in this systematic error stems from the fact that it constitutes the largest uncertainty of the combination. Notice that the b-JES systematic uncertainty, modeling the flavor response of b-JES, is a two-point systematic, i.e., based on the difference obtained with two methods, and thus falls into the category of systematic sources that could plausibly be uncertain. The confidence interval remains stable if any other of the systematic uncertainties considered are taken as themselves uncertain. A negligible bias in the confidence interval size of the b tagging and QCD radiation uncertainties is present due to the regularization of the correspondent correlation matrices.
@@ 5.2 Sensitivity to outliers
Figure 4: The plot shows the variation of the $68.3\%$ confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$ when one fictitious measurement is included in the combination. Each line represents the change in the confidence interval when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
@@
The conclusions of the last section would be significantly impacted if any of the combination inputs exhibited tension with the rest of the measurements. It is not uncommon to have input values that exhibit a significant tension and it could possibly happen for an updated LHC-Tevatron combination, or for a LHC Run $2$ combination that includes the ATLAS top mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 . However, since extending the combination to include additional measurements would require knowledge of the correlations between them and those listed in Tabs. 1 and 2, we introduce a fictitious measurement into the dataset to illustrate the properties of the GVM in such a scenario. Specifically, we consider adding a measurement with a central value of $m_{t}^{\mathrm{NEW}}=174.5$ GeV, a statistical uncertainty of $0.4$ GeV, and a global systematic uncertainty of $0.5$ GeV. We assume that the systematic uncertainty of the new measurement is uncorrelated with the uncertainties of the other measurements.
@@
If the fictitious measurement is added to the combination, without considering errors-on-errors, the result obtained using the BLUE approach is $m_{t}^{(1)}=172.91\pm 0.29\,\text{GeV}$. As anticipated, the inclusion of the new measurement shifts the central value of the combination towards higher values and results in a reduction of the confidence interval.
@@
Figure 3 illustrates the variations in the central value of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. On top of the systematics considered in the last section, the NEW systematic uncertainty on the fictitious measurement is also considered. In contrast to observations made in the previous section, the inclusion of the outlier in the dataset makes the effect of errors-on-errors non-negligible. The central value significantly shifts when the NEW systematic uncertainty is considered uncertain, moving back to that of the original combination. This is an important property of the GVM: while typically the presence of an outlier would significantly pull the result of a combination, within the GVM framework as the error-on-error parameter increases, the outlier is assigned lesser weight within the combination. Consequently, the central value shifts back towards that of the original combination, thereby diminishing the outlier’s effect.
@@
Similarly, Fig. 4 displays the variation in the half-size of the $68.3\%$ confidence interval as a function of $\varepsilon_{s}$. The plot shows how the inclusion of the fictitious measurement affects the combination. It demonstrates that when the NEW systematic uncertainty includes an associated error-on-error, the confidence interval can increase significantly, by up to 25%. This is the second relevant property of the GVM: The size of the confidence interval is sensitive to the internal consistency of the input data. Specifically, the less compatible the input measurements are, the more the confidence interval is subject to inflation when errors-on-errors are considered. This is because the GVM treats the tension in the dataset as an additional source of uncertainty resulting in an inflated confidence interval.
@@ 6 Conclusions
The Gamma Variance Model (GVM) provides a powerful statistical framework for addressing uncertainties in the assignment of systematic errors, informally referred to as errors-on-errors. Going beyond previous publications bib:Cowan2019 ; bib:Canonero2023 , we have extended the original model to include non-trivial correlations (not only 00 or ±$\pm 1$) and have provided formulas that greatly simplify its application in practice. As a by-product we have derived a useful connection between the BLUE method for combining results when there are non-trivial correlations and the corresponding likelihood method using nuisance parameters.
@@ Appendix A Profiling over NPs modeling correlated auxiliary measurements
The BLUE log-likelihood, with the covariance matrix defined as in Eq.(8), can be derived by profiling Eq.(10) over all the NPs $\theta_{s}^{i}$ when the auxiliary measurements $u_{s}^{i}$ are set to 00. This process involves profiling the likelihood recursively over all $\theta_{s}^{i}$ for a given $s$. Here, we demonstrate how to do this for $s=M$. To simplify the notation we define
@@ B.1 Profiled values of nuisance parameters
The method outlined in Sec. 4.1 for profiling the NPs can be extended to solve the score equations perturbatively for the GVM with arbitrary correlations, as defined in Eq. (16). Here we define the perturbative series for $\hat{\hat{\theta}}_{f(s,i)}$ as
$\hat{\hat{\theta}}_{f(s,i)}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\varepsilon_{s}^{4}\hat{\hat{\theta}}_{f(s,i)}^{(2)}+\cdots\,,$ (49)
where we introduce the composite index $f(s,i)=N(s-1)+i$ to denote $\theta_{s}^{i}$ for notational convenience. The first term in this series is given by
@@
This first-order result is also the solution to the quadratic likelihood given by Eq. (10). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative term $\varepsilon_{s}^{2n}\,\hat{\hat{\theta}}_{f(s,i)}^{(n)}$ is given by
@@
while $T_{f(s,k),n}$ and $S_{s,n}^{2}$ are defined as
$T_{f(s,k),n}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\cdots+\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(2n)}$ (54)
@@ B.2 Confidence intervals
Here, $\widehat{\sigma_{u_{s}}^{2}}$ is the MLE of $\sigma_{u_{s}}^{2}$, which can be computed by evaluating Eq. (15) at $\hat{\theta}_{s}^{i}$. The matrices $j^{-1}$ and $\tilde{j}^{-1}$ are defined by Eqs. (32) and (33), respectively, and are applied to the quadratic likelihood given in Eq. (10). Additionally, $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ and $\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ are submatrices of $j^{-1}$ and $\tilde{j}^{-1}$ that correspond to a fixed index $s$. Both $j^{-1}$ and $\tilde{j}^{-1}$ are functions of $\sigma_{u_{s}}^{2}$ and must be evaluated at $\widehat{\sigma_{u_{s}}^{2}}$. To simplify the computation of the Bartlett correction, the double-indexed NPs $\theta_{s}^{i}$ can be flattened into single-indexed variables by defining $f(s,i)=N(s-1)+i$, so that $\theta_{s}^{i}=\theta_{f(s,i)}$.
@@ Appendix C Generating toy data
where $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ is computed as in Eq. (58). These unbiased estimators were computed following the approach described in bib:Cox1968 ; bib:Codeiro1991 ; bib:Codeiro2014 . They are then used to generate auxiliary measurements, as well as the stochastic variables $v_{s}$, which follow a Gamma distribution:

### v4

@@ 5.2 Results
Figure 1 illustrates the variation in the central value (i.e., Maximum Likelihood Estimator) of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. Specifically, the plot shows the dependence of the central value $\hat{\mu}$ (here $\mu=m_{t}$) on the error-on-error parameter $\varepsilon_{s}$ for each systematic uncertainty. The nomenclature for the systematic components utilized here is the same used in bib:ATLAS/CMS , where a detailed explanation of what they describe can be found. The most important conclusion here is that the central value of the combination is quite robust to the presence of uncertain systematic errors. Specifically, the change in the central value remains always within $0.1$ GeV for the explored range of $\varepsilon_{s}$ values, which is well within the confidence interval of approximately $0.3$ GeV. Ultimately, all central values converge to the BLUE result as $\varepsilon_{s}$ approaches zero, meaning that the way we implement correlations in our model, as described in Sec. 2, is consistent. The reason why the b tagging and QCD radiation lines do not converge to the same point is due to the regularization of their associated correlation matrices, as previously described.
@@
Figure 2 confirms our previous conclusion: the combination is generally robust to the presence of uncertain systematic errors. However, an increase of about $10\%$ in the confidence interval is observed when the b-JES systematic uncertainty has a relative error of approximately $50\%$. The combination’s heightened sensitivity to uncertainties in this systematic error stems from the fact that it constitutes the largest uncertainty of the combination. Notice that the b-JES systematic uncertainty, modeling the flavor response of b-jets, is a two-point systematic, i.e., based on the difference obtained with two methods, and thus falls into the category of systematic sources that could plausibly be uncertain. The confidence interval remains stable if any other of the systematics considered are taken as themselves uncertain. A negligible bias in the size of the confidence interval stemming from the b tagging and QCD radiation uncertainties is present due to the regularization of the correspondent correlation matrices.
@@ 5.3 Sensitivity to outliers
Figure 4: The plot shows the variation of the $68.3\%$ half-size confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$ when one fictitious measurement is included in the combination. Each line represents the change in the confidence interval when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
@@
The conclusions of the last section would be significantly impacted if any of the combination inputs exhibited tension with the rest of the measurements. It is not uncommon to have input values that exhibit a significant tension and it could possibly happen for an updated LHC-Tevatron combination, or for a LHC Run $2$ combination that includes the ATLAS top-mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 . However, since extending the combination to include additional measurements would require knowledge of the correlations between them and those listed in Tabs. 1 and 2, we introduce a fictitious measurement into the dataset to illustrate the properties of the GVM in such a scenario. Specifically, we consider adding a measurement with a central value of $m_{t}^{\mathrm{NEW}}=174.5$ GeV, a statistical uncertainty of $0.4$ GeV, and a global systematic uncertainty of $0.5$ GeV. We assume that the systematic uncertainty of the new measurement is uncorrelated with the uncertainties of the other measurements.
@@
If the fictitious measurement is added to the combination, without considering errors-on-errors, the result obtained using the BLUE approach is $m_{t}=172.91\pm 0.29\,\text{GeV}$. As anticipated, the inclusion of the new measurement shifts the central value of the combination towards higher values and results in a reduction of the confidence interval.
@@
Figure 3 illustrates the variations in the central value of the combination when one of the systematic uncertainties listed in the legend is itself considered uncertain. On top of the systematics considered in the last section, the NEW systematic uncertainty of the fictitious measurement is also treated as potentially uncertain. In contrast to observations made in the previous section, the inclusion of the outlier in the dataset makes the effect of errors-on-errors non-negligible. The central value significantly shifts when the NEW systematic uncertainty is considered uncertain, moving back to that of the original combination. This is an important property of the GVM: while typically the presence of an outlier would significantly pull the result of a combination, within the GVM framework as the error-on-error parameter increases, the outlier is assigned lesser weight within the combination. Consequently, the central value shifts back towards that of the original combination, thereby diminishing the outlier’s effect.
@@
Similarly, Fig. 4 displays the variation in the half-size of the $68.3\%$ confidence interval as a function of $\varepsilon_{s}$. The plot shows how the inclusion of the fictitious measurement affects the combination. It demonstrates that when the NEW systematic uncertainty includes an associated error-on-error, the confidence interval can increase significantly, by up to 25%. This is the second relevant property of the GVM: The size of the confidence interval is sensitive to the internal consistency of the input data. Specifically, the less compatible the input measurements are, the more the confidence interval inflates when errors-on-errors are considered. This is because the GVM treats the tension in the dataset as an additional source of uncertainty resulting in an increase in the size of the confidence interval.
@@ 6 Conclusions
The Gamma Variance Model (GVM) provides a powerful statistical framework for addressing uncertainties in the assignment of systematic errors, informally referred to as errors-on-errors. Going beyond previous publications bib:Cowan2019 ; bib:Canonero2023 , we have extended the original model to include non-trivial correlations (not only 00 or ±$\pm 1$) and have provided formulas that simplify its application in practice by avoiding the need for numerical methods. We also derived a useful connection between the BLUE method for combining results under non-trivial correlations and the corresponding likelihood method using nuisance parameters. This connection was employed here to implement errors-on-errors when systematic uncertainties induce non-trivial correlations, but it can also be applied in other contexts.
We argue that the assumption of independent gamma distributions for estimates of variances is a plausible model that arises, e.g., if nuisance parameters are estimated through an average of Gaussian-distributed measurements. In a more general scenario, this assumption does not necessarily hold but the model should nevertheless provide more realistic inference than if one were to treat assigned variances as exactly known. The results one might obtain if the variance estimates followed some other bell-shaped distribution are expected to be qualitatively similar but dependent in general on the specific problem. Thus the precision of the inferences derived from the GVM is limited, but the model remains useful for identifying trends and understanding how an analysis would shift when uncertainties in the assignment of systematic errors are substantial.
@@ Appendix A Profiling over NPs modeling correlated auxiliary measurements
The BLUE log-likelihood of Eq. (4), with the covariance matrix defined as in Eq.(8), can be derived by profiling Eq.(10) over all the NPs $\theta_{s}^{i}$ when the auxiliary measurements $u_{s}^{i}$ are set to 00. This process involves profiling the likelihood recursively over all $\theta_{s}^{i}$ for a given $s$. Here, we demonstrate how to do this for $s=M$. To simplify the notation we define
@@ B.1 Profiled values of nuisance parameters
The method outlined in Sec. 4.1 for profiling the NPs can be extended to solve the score equations perturbatively for the GVM with arbitrary correlations, as defined in Eq. (18). Our aim is to solve
$\small\begin{split}&\sum_{j=1}^{N}\Gamma_{i}^{s}V_{ij}^{-1}\left(y_{j}-\mu-\sum_{p=1}^{M}\Gamma_{j}^{p}\hat{\hat{\theta}}_{f(p,j)}\right)\\ &+\frac{(1+2N\varepsilon_{s}^{2})\sum_{j=1}^{N}\left(\rho^{(s)}_{ij}\right)^{-1}\left(u_{s}^{j}-\hat{\hat{\theta}}_{f(s,j)}\right)}{v_{s}+2\varepsilon_{s}^{2}\sum_{i,j=1}^{N}\left(u_{s}^{i}-\hat{\hat{\theta}}_{f(s,i)}\right)\left(\rho^{(s)}_{ij}\right)^{-1}\left(u_{s}^{j}-\hat{\hat{\theta}}_{f(s,j)}\right)}=0\,,\end{split}$ (52)
where we introduce the composite index $f(s,i)=N(s-1)+i$ to denote $\theta_{s}^{i}$ for notational convenience. To solve the score equation, we define the perturbative series for $\hat{\hat{\theta}}_{f(s,i)}$ as
$\hat{\hat{\theta}}_{f(s,i)}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\varepsilon_{s}^{4}\hat{\hat{\theta}}_{f(s,i)}^{(2)}+\cdots\,.$ (53)
The first term in this series is given by
@@
This first-order result is also the solution to the quadratic likelihood given by Eq. (10). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative term $\varepsilon_{s}^{2n}\,\hat{\hat{\theta}}_{f(s,i)}^{(n)}$ is
@@
while $T_{f(s,i),n}$ and $S_{s,n}^{2}$ are defined as
$T_{f(s,i),n}=\hat{\hat{\theta}}_{f(s,i)}^{(0)}+\varepsilon_{s}^{2}\hat{\hat{\theta}}_{f(s,i)}^{(1)}+\cdots+\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(2n)}\,,$ (58)
@@
When variations of this model are employed, the same perturbative procedure must be used. To solve the perturbative equation for $\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(n)}$, one substitutes $\hat{\hat{\theta}}_{f(s,i)}$ with $T_{f(s,i),n}$ throughout the score equation, except in the denominator where $T_{f(s,i),n-1}$ is substituted instead (due to the presence of the $\varepsilon_{s}^{2}$ multiplicative factor). Since all perturbative terms in $T_{f(s,i),n-1}$ have already been determined recursively, one can isolate $\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{f(s,i)}^{(n)}$ and solve for it directly.
@@ B.2 Confidence intervals
Here, $\widehat{\sigma_{u_{s}}^{2}}$ is the MLE of $\sigma_{u_{s}}^{2}$, which can be computed by evaluating Eq. (17) at $\hat{\theta}_{s}^{i}$. The matrices $j^{-1}$ and $\tilde{j}^{-1}$ are defined by Eqs. (34) and (35), respectively, and are applied to the quadratic likelihood given in Eq. (10). Additionally, $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ and $\tilde{j}^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ are submatrices of $j^{-1}$ and $\tilde{j}^{-1}$ that correspond to a fixed index $s$. Both $j^{-1}$ and $\tilde{j}^{-1}$ are functions of $\sigma_{u_{s}}^{2}$ and must be evaluated at $\widehat{\sigma_{u_{s}}^{2}}$. To simplify the computation of the Bartlett correction, the double-indexed NP $\theta_{s}^{i}$ can be flattened into single-indexed variable by defining $f(s,i)=N(s-1)+i$, so that $\theta_{s}^{i}=\theta_{f(s,i)}$.
For variations of the log-likelihood definition, such as the one in Eq. (41), the Bartlett factors retain the same functional form. One need only apply the matrices $j^{-1}$ and $\tilde{j}^{-1}$ to the corresponding quadratic log-likelihood, which, for instance, arises from Eq. (41) in the limit $\varepsilon_{s}\to 0$.
@@ B.3 Goodness-of-fit
If variations of the likelihood are employed, note that instead of $N$, the dimension of the NP basis must be used. For instance, in the case of the log-likelihood in Eq. (41), one would substitute $(2M_{s}+M_{s}^{2})$ in place of $(2N+N^{2})$.
@@ Appendix C Generating toy data
where $j^{\boldsymbol{\theta}_{s}\boldsymbol{\theta}_{s}}$ is computed as in Eq. (62). These unbiased estimators were computed following the approach described in bib:Cox1968 ; bib:Codeiro1991 ; bib:Codeiro2014 . They are then used to generate auxiliary measurements $u_{s}^{i}$, as well as the stochastic variables $v_{s}$, which follow a Gamma distribution: